how do you factor 4x^2-13x-12?
Found 2 solutions by stanbon, TakeATuition.com:
Answer by stanbon(75887)

factor 4x^2-13x-12
= 4x^2 -16x+3x-12
= 4x(x-4)+3(x-4)
= (x-4)(4x+3)
==================
Cheers,
Stan H.

Answer by TakeATuition.com(57)

SOLUTION -
4X^2 - 13X - 12 = 0
4X^2 - 16X + 3X - 12 = 0
4X(X - 4) + 3(X - 4) = 0
(X - 4)(4X + 3) = 0
X - 4 = 0
X = 4
4X + 3 = 0
4X = -3
x = -3/4
FOR MORE HELP VISIT :www.TakeATuition.com